OpenStudy (anonymous):

Find the points of inflection for the curve; y = x(x² -5)² help please.. thank you.

OpenStudy (mimi_x3):

Find the derivative of x(x^2-5)^2 dy/dx = 5(x^4-6x^2+5) Then find the second derivative f'' = dy^2/dx^2 = 20x(x^3-3) 20x(x^3-3)=0 , then find x and y that is the poi

OpenStudy (anonymous):

double differenciate and put y'' = 0.